OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (kropot72):

Sustitute the given values: C = (0.765 * 100) + 0.06(0.765 * 100) = 76.5 + (0.06 * 76.5) = 76.5 + 4.59 = $81.09

OpenStudy (anonymous):

You just plug in 100 for x C = 0.765(100) + 0.06(.0765*100) Answer is $81.09